HUBBLE SPACE TELESCOPE DAILY REPORT #2852 PERIOD COVERED: 0000Z (UTC) 04/16/01 - 0000Z (UTC) 04/17/01 Daily Status Report as of 107/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Five Sets of WF/PC-2 8827 (Cycle 9 Supplemental Darks pt2/3) The WF/PC-2 was used to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The proposal completed nominally. 1.2 Completed Two Sets of STIS/CCD 8864 (CCD Dark Monitor-Part 2) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks for the CCD. The proposal completed nominally. 1.3 Completed STIS/CCD/MA1 8606 (Determining the Nature of the Variable Absorption in AGN: Monitoring NGC 3783 with HST And Chandra) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to investigate Seyfert 1 galaxies, many of which show intrinsic UV absorption lines, characterized by high moderate widths, significant outflow velocities, and variability on time scales as small as days. Seyferts with UV absorption also show variable X-ray ``warm absorbers'', characterized by O VII and O VIII absorption edges, which suggests a common origin. Variability monitoring is the key to understanding the absorbers, by providing their radial locations, densities, and evolution in ionization, column density, velocity, and coverage of the inner active nucleus. The proposal completed with no problems. 1.4 Completed Nine Sets of WF/PC-2 9244 (POMS Test Proposal: WFII Parallel Archive Proposal Continuation) The WF/PC-2 was used to perform a generic target version of the Archival Pure Parallel program. The program will be used to take parallel images of random areas of the sky, following the recommendations of the Parallels Working Group. The observations completed with no anomalous activity. 1.5 Completed Two Sets of WF/PC-2 8812 (Cycle 9 Internal Monitor) The WF/PC-2 was used to obtain routine internal monitors for WFPC2 in order to monitor the health of the cameras. A variety of internal exposures are obtained in order to provide a monitor of the integrity of the CCD camera electronics in both bays {gain 7 and gain 15}, a test for quantum efficiency in the CCDs, and a monitor for possible buildup of contaminants on the CCD windows. There were no reported problems. 1.6 Completed STIS/CCD 8865 (Bias Monitor-Part 2) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4, to build up high-S/N superbiases and track the evolution of hot columns. The proposal completed with no anomalous activity. 1.7 Completed Two Sets of WF/PC-2 8683 (Imaging Of Brightest Cluster Galaxies: The High End Of The Black Hole Mass Distribution) The WF/PC-2 was used to make kinematic black hole detections in galaxies to decide whether they indicate that the mass correlates with both optical luminosity and radio power. The observation completed with no reported problems. 1.8 Completed Four Sets of STIS/CCD 9248 (Probing the Large Scale Structure: Cosmic Shear Observations) The Space Telescope Imaging Spectrograph (CCD) was used to measure the distortion of light bundles from distant galaxies that probe the statistical properties of the intervening inhomogeneous {dark} matter distribution. Its tidal gravitational field distorts the observable image shapes thereby causing a coherent ellipticity pattern {Cosmic Shear}. The observations completed nominally. 1.9 Completed Two Sets of STIS/CCD 8840 (Read Noise Monitor) The Space Telescope Imaging Spectrograph (CCD) was used to measure the read noise of all the amplifiers on the STIS CCD using pairs of bias frames. Full frame and binned observations are made in both Gain 1 and Gain 4, with binning factors of 1 x 1, 1 x 2, 2 x 1 and 2 x 2. All exposures are internals. The observations completed with no reported problems. 1.10 Completed WF/PC-2 8581 (A Search For Low-Mass Companions To Ultracool Dwarfs) The WF/PC-2 was used to search for very low-mass {VLM} companions to a complete sample of 120 late-M and L dwarfs, drawn mainly from the 2MASS and SDSS surveys. The primary goal is to determine the multiplicity of M < 0.1 M_odot dwarfs. In particular, we aim to identify binary systems suitable for long-term astrometric monitoring and mass measurement, and systems with cool, sub-1000K companions. The proposal completed with no reported problems. 1.11 Completed Two Sets of STIS/CCD 8607 (Completing the Local AGN Inventory: The AGN Content of Composite Nuclei) The Space Telescope Imaging Spectrograph (CCD) was used to increase the knowledge of the local space density of AGNs which is of fundamental importance to a number of astrophysical problems. A significant fraction of nearby galaxies have nuclei whose spectra are intermediate between those of LINERs and nuclear starbursts. These ``transition objects'' may be composite systems that contain both a central AGN and circumnuclear star formation. We will test this hypothesis by obtaining spatially-resolved optical spectra of a well-defined sample of 15 nearby transition nuclei selected from the extensive ground-based survey of Ho, Filippenko, and Sargent. The observations completed with no reported problems. 1.12 Completed Four Sets of WF/PC-2 8815 (Cycle 9 Earth Flats) The WF/PC-2 was used to monitor flatfield stability. This proposal obtains sequences of Earth streak flats to construct high quality flat fields for the WF/PC-2 filter set. These flat fields will allow mapping of the OTA illumination pattern and will be used in conjunction with previous internal and external flats to generate new pipeline superflats. The proposal completed nominally. 1.13 Completed Two Sets of WF/PC-2 8824 (Decontaminations and Associated Observations Pt. 3/4) The WF/PC-2 was used for the monthly WFPC2 decons. Also included are instrument monitors tied to decons: photometric stability check, focus monitor, pre- and post-decon internals {bias, intflats, kspots, & darks}, UV throughput check, VISFLAT sweep, and internal UV flat check. There were no reported problems. 1.14 Completed S/C 5582 (FOC/48 Monthly Activation (Camera Section Only): Cycle 4) This proposal was executed as scheduled, and no anomalies were noted. 1.15 Completed FOC/96 7923 (FOC F/96 Turn-on) The Faint Object Camera (f/96) was used to take an internal flat and a dark to ensure that the 30-day limit for the FOC is not exceeded. The observations were executed as scheduled, and no anomalies were noted. 1.16 Completed FGS/1 8774 (Calibrating the Mass-Luminosity Relation at the End of the Main Sequence) The FGSs-1R was used to calibrate the mass-luminosity relation {MLR} for stars less massive than 0.2 Msun, with special emphasis on objects near the stellar/brown dwarf border. Our goals are to determine Mv values to 0.10 magnitude, masses to 5%, and more than double the number of objects with masses determined to be less than 0.20 Msun. The proposal completed with no reported problems. 1.17 Completed STIS/CCD/MA1 8712 (The Wind Accretion Wake In A Detached Binary System) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to observe RW Hya {= HD 117970}, the only known detached binary system, where direct observational evidence for wind accretion onto a white dwarf is found. UV observations will enable the investigator to advance knowledge in a field which has for long had to remain a theoretical subject. RW Hya is an eclipsing system containing a mass-losing M-giant and a hot white dwarf on circular orbits {370 days}. The observation completed with no reported problems. 1.18 Completed STIS/MA1 8843 (Cycle 9 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA1) was used to perform the routine monitoring of the MAMA detector dark noise. The proposal completed nominally. 1.19 Completed STIS/CCD/MA1 8665 (The Starburst - AGN Connection: The Nature of the UV-bright Core in NGC 4303) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to observe NGC-4303, to date the best example of the claimed starburst- AGN connection in active galaxies. The observations completed with no reported problems. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 8 Successful: 8 Scheduled Re-acquisitions: 8 Successful: 8 2.2 FHST Updates: Scheduled: 20 Successful: 20 2.3 Operations Notes: The SSR EDAC error counter was cleared three times per ROP SR-1A. A TTR was generated when data dropouts occurred, starting at 106/1527Z. The ephemeris table was uplinked at 107/0005Z as directed by ROP DF-07A. The STIS MCE-2 reset at 107/035110Z while the low voltage was on and while outside any SAA interval. The STIS flight software error counter was reset at 107/0403Z, using ROP NS-12. MAMA-2 will be recovered, via normal SMS commanding, at 1081740Z when the high voltage is next turned on. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /CAW